Suppose music producer Mackleless is considering signing the up and coming artist AriannePetite. The record sales of Mackleless have a market beta of 1, while Arianne Petite’s record sales have a market beta of 2. The risk-free rate is currently 5%, while the expected market return is 10%.

a. What is the relevant discount factor for Arianne Petite’s record sales?

Suppose Arianne Petite requires a sign-up bonus of $60 million, and the cash flows from her record sales are expected to be $10 million for the next 10 years.

b. Is signing Arianne Petite a good idea for Mackleless?

c. Does your answer change if Mackleless’ beta is used for calculation of the discount

factor?
Answer #1

Ri = Rf + B(Rm-Rf) where Rf = Risk free return, Rm = Return from market, B = Beta factor, Ri= individual return

Let 'a' be Mackleless and 'b' be Arianne Petite’

Ra = 5 + 1(10-5) = 10%

Rb = 5 + 2(10-5) = 15%

a) relevant discount factor for Arianne Petite’s record sales is 15%.

The cash flows from her record sales are expected to be $10 million for the next 10 years. So discount rate of 15% for 10 years will be 5.0188. So, NPV = (5.0188*10) - 60 = -9.812 million.

b) It is NOT a good idea for Mackleless to sign as it generates negative NPV.

c) Discount rate of 10% for 10 years will be 6.145. So, NPV =(6.145*10) - 60 = +1.45. Since the NPV is positive, its advised to go ahead with the signing. So the answer will obviously change.

Please comment in case of any query regarding the solution.